mancity:

According to the 1998 Information Please Almanac, the official speed record for running one mile was held by Richard Webster of England in 1865, with a time of 4 minutes 36.5 seconds. Roger Bannister, also of England, was the first to run a mile in less than four minutes, clocking a time of 3 minutes 59.4 seconds in 1954. There are (of course) 60 seconds in each minute. Therefore, we could convert both times into seconds, writing them as follows: a1865 = 276.5 seconds and a1954 = 239.4 seconds. Predict the world record for the mile in the year 2010, leaving your answer in seconds.

Vocaloid:

Judging by the notation, it wants you to treat the times as an arithmetic sequence an = a1 + (n-1)d create two separate equations for n = 1865 and n = 1954 a1865 = a1 + (1865-1)d = 276.5 a1954 = a1 + (1954-1)d = 239.4 if you subtract the two equations, you can eliminate a1 and solve for d then you can go back to either equation, plug in d, evaluate a1 finally, from there, you can use the original equation, plug in a1 and n = 2010, d, and solve for an